The chlorofluorocarbon refrigerant Freon-12, CCl2F2, was banned in 1995 to help protect the ozone layer. Most new air conditioning units now use the re- frigerant tetrafluoroethane, CF3CH2F. The enthalpy of vaporization of Freon-12 is 155 Jg–1 and that of tet- rafluoroethane is 215.9 Jg–1. Estimate the number of grams of Freon-12 that must be vaporized to freeze a tray of 16 one-ounce (1 oz = 28 g) ice cubes with the water initially at 18°C. How many grams of tetrafluo- roethane are required to perform the same task?

Answer:1180 g, 848 g
Answer #1

Amount of heat need to be removed(q) = m*s*DT + n*DHfus

m = mass of water to be frozen = 16 oz*(28 g/1 oz) = 448 g

s = specific heat of water = 448 j/g.c

DT = 18-0

n = no of mole of water to be frozen = 448/18 = 24.9 mole

DHfus = 6.01 kj/mol

q = 448*4.184*(18-0)+24.9*6.01*10^3

q = 1.834*10^5 joule

q = 1.834*10^2 kj

Amount of Freon-12 that must be vaporized(m1) = q/C

q = 1.834*10^5 joule , C = heat capacity of Freon-12 = 155 j/g

m1 = 1.834*10^5/155 = 1183.2 g

Amount of tet- rafluoroethane that must be vaporized(m2) = q/C

q = 1.834*10^5 joule , C = heat capacity of tet- rafluoroethane = 215.9 j/g

m2 = 1.834*10^5/215.9 = 849.5 g